OpenStudy (anonymous):

Can someone give me 5 examples of multiplicative inverses please

OpenStudy (anonymous):

A multicative inverse is when you can multiply a number by something to get 1 so an example would be what can we multiply 8 by to get 1 \[8 \times ? = 1\] Since we should know multiplying by a reciprocal gets 1 then \[8 \times \frac{ 1 }{ 8 } = 1\] This is what the multiplicative identity is so four other examples could be \[\frac{ 2 }{ 3 } \times \frac{ 3 }{ 2 } = 1\] \[1 \times \frac{ 1 }{ 1 } = 1\] \[\frac{ 13 }{ 100 } \times \frac{ 100 }{ 13 } = 1\] \[\frac{ 6789 }{ 14 } \times \frac{ 14 }{ 6789 }=1\]